मराठी

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).

Advertisements
Advertisements

प्रश्न

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).

बेरीज
Advertisements

उत्तर

P(A) = 0.42, P(B) = 0.48

P(A and B) = P(A ∩ B) = 0.16

P(not A) = P(A’) = 1 – P(A)

= 1 – 0.42

= 0.58

shaalaa.com
Concept of Probability - Probability of 'Not', 'And' and 'Or' Events
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 14: Probability - EXERCISE 14.2 [पृष्ठ ३०७]

APPEARS IN

एनसीईआरटी Mathematics [English] Class 11
पाठ 14 Probability
EXERCISE 14.2 | Q 17. (i) | पृष्ठ ३०७

संबंधित प्रश्‍न

If `2/11` is the probability of an event, what is the probability of the event ‘not A’.


A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is

  1. a vowel
  2. an consonant

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(A or B).


The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?


In a class of 60 students, 30 opted for NCC, 32 opted for NSS and 24 opted for both NCC and NSS. If one of these students is selected at random, find the probability that

  1. The student opted for NCC or NSS.
  2. The student has opted neither NCC nor NSS.
  3. The student has opted NSS but not NCC.

A dice is thrown. Find the probability of getting a prime number


A dice is thrown. Find the probability of getting a multiple of 2 or 3.

 

In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers


In a simultaneous throw of a pair of dice, find the probability of getting a doublet of odd numbers


In a simultaneous throw of a pair of dice, find the probability of getting an even number on one and a multiple of 3 on the other


In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 6


In a simultaneous throw of a pair of dice, find the probability of getting a sum less than 7


In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 7


In a simultaneous throw of a pair of dice, find the probability of getting a number greater than 4 on each die


In a simultaneous throw of a pair of dice, find the probability of getting a total of 9 or 11


In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.

 

Two dice are thrown. Find the odds in favour of getting the sum 4.


Two dice are thrown. Find the odds in favour of getting the sum 5.

 

 


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white .


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white and odd numbered .


Fill in the blank in the table:

P (A) P (B) P (A ∩ B) P(A∪ B)
\[\frac{1}{3}\] \[\frac{1}{5}\] \[\frac{1}{15}\] ......

A card is drawn at random from a well-shuffled deck of 52 cards. Find the probability of its being a spade or a king.


One of the two events must occur. If the chance of one is 2/3 of the other, then odds in favour of the other are


The probability that a leap year will have 53 Fridays or 53 Saturdays is


If the probability of A to fail in an examination is \[\frac{1}{5}\]  and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is

 
 

A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is


Two dice are thrown simultaneously. The probability of getting a pair of aces is


An urn contains 9 balls two of which are red, three blue and four black. Three balls are drawn at random. The probability that they are of the same colour is


A box contains 6 nails and 10 nuts. Half of the nails and half of the nuts are rusted. If one item is chosen at random, the probability that it is rusted or is a nail is


Three numbers are chosen from 1 to 20. The probability that they are not consecutive is


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×